### v2

# Diversity and stylization of the contemporary user-generated visual arts in the complexity-entropy plane

Abstract
The advent of computational and numerical methods in recent times has provided new avenues for analyzing art historiographical narratives and tracing the evolution of art styles therein. Here, we investigate an evolutionary process underpinning the emergence and stylization of contemporary user-generated visual art styles using the complexity-entropy (C-H) plane, which quantifies local structures in paintings. Informatizing 149,780 images curated in DeviantArt and Bēhance platforms from 2010 to 2020, we analyze the relationship between local information of the C-H space and multi-level image features generated by a deep neural network and a feature extraction algorithm. The results reveal significant statistical relationships between the C-H information of visual artistic styles and the dissimilarities of the multi-level image features over time within groups of artworks. By disclosing a particular C-H region where the diversity of image representations is noticeably manifested, our analyses reveal an empirical condition of emerging styles that are both novel in the C-H plane and characterized by greater stylistic diversity. Our research shows that visual art analyses combined with physics-inspired methodologies and machine learning, can provide macroscopic insights into quantitatively mapping relevant characteristics of an evolutionary process underpinning the creative stylization of uncharted visual arts of given groups and time.
Keywords creativity |||| diversity |||| image representation |||| stylization |||| complexity-entropy plane
1 Introduction
The contemporary visual arts scene embraces a wide array of artistic forms and creative standards that are gaining popularity through information-sharing platforms. While there have been significant historical examinations of how creative visual art styles have evolved, the quantitative analysis of creative stylization is a relatively recent development. We introduce a research design that aims to document the characteristics of the emergence of artistic movements and their subsequent stylizations. By testing more refined empirical data, we improve upon the previous statistical physics approach. Our interdisciplinary research design incorporates two distinct fields.
First, theoretical definitions of creativity have identified two sequential phases in the recognition of creativity. In the first phase, creativity is seen as “the intentional arrangement of cultural and material elements in a way that is unexpected for a given audience [1].” In the following phase, the creativity finds its “relevance in the new context,” where it contributes to the solution of a problem it assiduously formulated [2]. While the first phase that “caused scandal” may not necessarily result in the second phase, it often serves as a catalyst for the development of novel artistic styles. Art movements like Surrealism and Dadaism, among others, have used this disruption to challenge conventional norms and spark creative revolutions [3].
Upon the concatenation of the initial disruption and the emergence of a new style, a transformative process takes place. This process influences the creative acts of subsequent artists through the observation, analysis, and internalization of the new style. By assimilating and building upon these innovations, artists contribute to the dynamic evolution of art [4, 5]. Therefore, to gain a comprehensive understanding of the creative process, it is essential to consider the two phases separately. By analyzing the initial novelty and the emergence of a new style as distinct phenomena, we can better appreciate the individualized contributions of each artwork as an original mutation. This analytical approach allows for a more nuanced interpretation of artistic development within an enhanced spatiotemporal framework.
Second, while theoretical studies of creativity in the domain of art can be traced back to earlier eras, contemporary efforts to quantitatively assess and measure relevant characteristics have only recently gained momentum [6, 7]. Through the development of computational and robust methods for analyzing visual arts and the accumulation of large-scale digital art images, scientists have characterized and uncovered various interesting features within a broad range of artworks. Computational studies of visual art have provided both synchronic and diachronic insights into zeitgeist and evolution of visual arts. Previous studies of visual art styles have been advanced with statistical data of spatial parameters, including chromatic properties [8], fractals [9], wavelets [10], compression ensemble approach [11], and concepts of entropy [12, 13, 14].
In particular, complexity (C)-entropy (H) plane as a statistical tool has proven to intuitively map a diachronic evolution of art-historical styles [13, 15]. In statistical physics, measurements of C and H based on the distribution of local ordering patterns were initially used to characterize time-series signals; these measurements have since been generalized to analyze patterns of 2D images [16]. C represents the degree to which the local order patterns deviate from both a random and homogeneous distribution, and H represents the degree of disorder in the image’s pixel organization. C reflects the degree to which the objects within an image are spatially bounded or interrelated (e.g., Impressionist paintings tend to have low C due to the use of different types of brush patterns; different local patterns are detected uniformly on a painting), whereas H reflects the degree to which the objects of an image are more clearly outlined or exhibit fluidity between them (e.g., Mondrian’s minimal geometric paintings tend to have low H) (refer to Methods; complexity-entropy (C-H) measures of visual artwork images).
Sigaki et al.’s approach [13] appraised the localization of the C and H of 137,364 visual artwork images from WikiArt, and gained a view of the hierarchical clustering structure of 92 art-historical styles (between Renaissance and Contemporary/Postmodern Art) in the C-H plane. The study regarded the dynamics of transitions of different artistic styles in the C-H plane as evolutionary. This suggests that decoding the latent dynamics that have driven the path of the styles in the C-H plane would reveal the evolutionary process underpinning the emergence and stylization of a specific group of visual arts in terms of the two aforementioned phases of creativity and stylization. Yet, the deterministic properties of the localized information of artworks in the bounded C-H space raises a question of how to estimate an evidence of creative stylization of visual arts in the C-H plane.
Here, we consider the indication of diverse image representations of a specific timeframe and groups (i.e., intragroup image diversity) as a conditional phase of the visual artistic stylization in the C-H plane. In this context, we conduct an exploratory data analysis (EDA) [17] on stylization of 149,780 visual art images from quasi-canonical “user-generated arts [18]” of DeviantArt and Bēhance platforms of a given timeframe (2010-2020) in the C-H plane, as an extension of the temporal evolution of art styles by Sigaki et al. Through empirically identifying the relationships between the intragroup image diversity and the corresponding local information of the C-H space, we hypothesize that there will be a particular C-H movement (i.e., mutations) of a specific group of visual artworks over a given timeframe for their stylistic diversity to manifest in the C-H plane. To test this hypothesis, we ask the following research questions.
RQ 1. Can the C-H plane, which effectively characterized the styles of art-historical paintings, also capture the temporal stylistic evolution (i.e., C-H trajectories) of contemporary user-generated visual arts? If so, what would these trajectories represent?
RQ 2. How is the average C-H position of user-generated visual arts at a given time related to the intragroup image diversity? Can we comprehend the relationship between C-H positions and the intragroup image diversity through the local information of the C-H space, specifically focusing on the diversity of image representations that can be expressed in particular regions?
To address the issues in the aforementioned questions, we conduct the following analyses.
1) Through validating the applicability of our image set conveyed by the C-H measures and their robustness, we reveal characteristics of temporal stylistic transitions of quasi-canonical user-generated visual arts from DeviantArt and Bēhance in the C-H plane. Moreover, we explore sub-visual art fields within the user-generated visual arts that have significantly influenced the temporal stylistic transitions in the C-H plane.
2) We use two types of similarity measures - cosine similarity and Jaccard similarity, on two multiple image representation space – image embeddings through a pre-trained Residual Network (ResNet) architecture [19] and Scale-invariant Feature Transform (SIFT) features [20], to measure the image diversity/degree of dissimilarity in a specific C-H region. The low- (e.g., visual elements such as lines, contours, height, edges, angles, dots, colors, etc.) and high-level image features (e.g., themes of shapes and objects comprised of low-level features) extracted from the two methods aggregately encompass art style-level explicability of images [21].
3) In addition, we use the autoregressive moving-average (ARMA) model to statistically examine a temporal relationship between the average C-H positions of artworks within a given period and the average dissimilarities of their multi-level image features. Finally, we investigate the diversity of image representations that can be expressed in different areas of the C-H space and unveil an empirical condition for the emergence of styles that are not only novel in the C-H plane, but also characterized by greater stylistic diversity.
2 Results
2.1 Stylization of curated visual arts in the C-H plane: DeviantArt and Bēhance (2010-2020)
To account for the stylization of contemporary visual artworks mainly distributed via online communication channels of information, we investigate DeviantArt and Bēhance, serving as massive online visual art platforms involving various creative fields. Since there are huge collections of artworks in both platforms, we specifically focus on representative and quasi-canonical subsets of daily promoted user-generated visual artworks (“Daily Deviation” and “Best of Bēhance”) curated by those platforms. This allowed us to process manageable and significant amounts of data [22, 23]. The image set used for complexity-entropy (C-H) measurement is described in Table SI 1.
We map the yearly average C and H values of 149,780 user-generated visual artwork images from DeviantArt and Bēhance (2010 - 2020) (see Methods; Complexity-entropy (C-H) measures of visual artwork images, for a detailed explanation of the calculation of C and H values). For comparison, we also overlay the C-H values of conventional art historical paintings from the WikiArt dataset [24] as Sigaki et al.’s experiment [13]. The WikiArt dataset is composed of 26,415 paintings, primarily collected from Western art history, spanning the period of 1301 to 2016 CE. The art historical periods depicted in our projection of the evolution of visual arts are virtually identical to those identified by Sigaki et al.
Figure 1: Group-level temporal transitions of visual art styles in the C-H plane. (a) Contemporary user-generated visual art styles of DeviantArt and Bēhance (2010 - 2020) compared to the main divisions of art historical periods mapped in the C-H plane. The result reveals that the average C-H values of art historical styles have been shifted towards the higher C and lower H area over time. (b, c) Yearly C-H trajectory of contemporary user-generated visual arts from DeviantArt’s Daily Deviations and Bēhance’s the Best of Bēhance (refer to Data (Materials)). Each C-H plane is composed of multiple elliptical areas, with each ellipse representing 95% confidence interval bounds from the yearly average C-H values of the intragroup visual artworks. The multiple CI ellipses and the yearly C-H trajectories connecting them altogether describe yearly transitions of visual art styles within each platform. Also, insets are included to show cumulative number of samples over the given time frame.
Fig. 1 demonstrates that both platforms’ C-H trajectories are tended towards the upper left C-H region compared to the previous periods, although their average position shows a slight difference (refer to Table SI 2 for more information regarding the difference in the overall positions of the two platforms in the C-H plane).
Also, a closer look on the C-H values of DeviantArt and Bēhance reveals that the averaged position of the two platforms have varied significantly over time. We cross-validated the temporal difference in C-H values on each platform by evaluating the predictive accuracies of our dataset’s C-H values using four machine learning algorithms. All the classification models predict the curation year of paintings with probabilities (DeviantArt: $\sim$15%, Bēhance: $\sim$13.7%) significantly higher than the chance level (Fig. SI 1).
The C-H trajectories of temporal transitions of DeviantArt and Bēhance visual art styles over a decade extend that of the transition between early 20th century and 1970s (e.g., from painterly/optic to linear/haptic). Observing the beginning and ending years (Fig. 1b, c), the C of DeviantArt’s visual art increased by 0.002 (from 0.123 to 0.125) while their H decreased by 0.005 (from 0.84 to 0.835). In the case of Bēhance, there was a rise in C by 0.01 (from 0.12 to 0.13) and a decline in H by 0.11 (from 0.81 to 0.7) over the same period. The group-level temporal transitions of visual art styles among the two platforms reveal indirect but eventual tendencies towards styles with higher C and lower H. Consequently, the chronological C-H trajectories of the contemporary user-generated visual arts of both platforms from 2010 to 2020 indicate that their stylizations altered in a macroscopically similar manner, despite the differences in direction details within each platform.
Meanwhile, the yearly average C-H values of Bēhance are notably distinct from those of DeviantArt and the major art historical periods. We explore which sub-visual art fields substantially influenced the C-H positions of Bēhance over time, by separately examining the temporal C-H movements of five creative fields that account for the greatest proportion of samples in Bēhance dataset. We observe the yearly average C-H values of samples from Illustration, Graphic Design, Character Design, Animation, and Fine Arts fields in Bēhance. Fig. SI 2 shows that all the five fields’ yearly average C-H values are specifically directed towards a similar C-H area (C: 0.1 - 0.3, H: 0.65 - 0.85) through a decade of movements. Overall, the C-H movements of the fields reveal a tendency for homogeneous transitions of visual art styles in Bēhance during the time period.
Extending Sigaki et al.’s analysis on the evolution of art styles allows us to discover the group-level temporal stylizations of quasi-canonical visual artworks in DeviantArt and Bēhance (2010 – 2020) aligning with that of the modern art identified by Sigaki et al. Overall, the C-H movements of the platforms are characterized by an increase in high-level visual structure [25] of artworks incorporating geometrical object-oriented patterns; the emergence and taking up of an artistic style with clearer and simpler visual elements.
Through structural notions of creativity and stylization, it is possible to infer the strengthened stylization of the contemporary user-generated visual arts towards a more linear/haptic region in the C-H plane. One possible interpretation is the “bijectively homomorphic” (i.e., isomorphic) [26] stylization of the artworks. In other words, random and diverse visual artistic representations of artworks interact and homogenize while linearly transforming underlying characteristics of art-makings over time. The significance of individual artists’ innovation being proportional to the extent of its influence on others [27], fuels a competitive and mimetic process under uncertainty, paradoxically resulting in a cycle of imitation as a necessary step toward creativity [28]. Also, in any given institution, the number of “insiders” on the curatorial committee is fewer than expected, resulting in “a durable and recognizable pattern of aesthetic choices [29]”. In response to the inherent “symbolic uncertainty” in the platform of visual arts, individual artists may be subject to “mimetic isomorphism,” or mimicking others’ distinguished artistry as the most cost-effective strategy for gaining insider acceptance [30]. One could argue that the constantly quickened dissemination of information and images, influenced and expected by visual content structures on the platforms, has accelerated the homogenization of contemporary user-generated visual art styles.
2.2 Spatiotemporal relationships between image diversity and the C-H space
We investigate how the average C-H positions of user-generated visual arts from DeviantArt and Bēhance moved over time. In this section, we attempt to reveal the spatiotemporal relationships between image diversity and the local information of the C-H space, so that we can understand the previous C-H movements over time in terms of intragroup diversity and stylization.
In order to assess the intragroup diversity of image representations from the two platforms in the C-H space, we use two different similarity measures on two types of image features. First, we use cosine similarity of the artwork’s image embeddings (IE): multi-level image representations comprising both low- (100-dimensional) and high-level (100-dimensional) image features. Further, we use the Jaccard similarity coefficient of low-level features (i.e., keypoints) of two images matched by the standard SIFT algorithm. Fig. 2 shows the detailed pipelines of the two image similarity measures. The analytical independence of the two similarity measures and the distance in the C-H plane verifies that their pairwise correlations are fairly weak (range: 0.003 - 0.07; see Fig. SI 3).
Figure 2: Image representations in the complexity-entropy (C-H) plane and pipelines of image similarity measures through different image representation spaces. (a) Three reputable exemplary masterpieces chosen and localized in the C-H plane according to the given embedding parameters (dx = dy = 2), to aid viewers’ understandings of the image representations the plane encompasses. (b) A schematic diagram of image similarity computation pipeline through a neural network. We measure pairwise cosine similarities of a 200-dimensional embedding vector including both low- and high-level feature maps extracted from individual visual artworks. (c) A schematic diagram of image similarity computation pipeline through SIFT descriptor. We measure pairwise Jaccard similarities of SIFT descriptor matching a 128-dimensional vector of low-level features extracted from individual visual artworks. In addition to the image representations the C-H plane encompasses, we adopt and use two different types of image processing methods to obtain multi-level image features and measure their similarities: ResNet architecture and SIFT algorithm. Through our approach, we investigate characteristics of image representations in the C-H plane with image similarity measures that aggregate both low- and high-level features. The implemented images (1 – 3.) are “Kompozicija II by Piet Mondrian, 1929 (WikiArt),” “Rhythm by Robert Delaunay, 1912 (WikiArt),” and “The starry night by Vincent van Gogh, 1889 (WikiArt).” All the used image samples are available in the public domain.
We first examine the relationship between the average C-H positions of user-generated visual arts from DeviantArt and Bēhance over the specified timeframe (2010-2020) and their intragroup image diversity. In doing so, we apply the ARMA model to statistically regress changes in average values from the two image similarity measures on the average values of the complexity C and entropy H over the given years.
Regressions Predicting Similarity with ARMA Errors (2010-2020)
Variables IE Similarity IE Similarity SIFT Similarity SIFT Similarity
(DeviantArt) (Bēhance) (DeviantArt) (Bēhance)
(P = 3, Q = 1) (P = 1, Q = 1)
Entropy H Mean -3.650*** -1.018*** -0.043*** -0.135*
(0.250) (0.177) (0.004) (0.055)
Entropy H Variance -5.842*** -1.632*** -0.087*** -0.278**
(0.770) (0.376) (0.010) (0.105)
Complexity C Mean -4.934*** -3.801*** -0.061*** -0.263+
(0.342) (0.400) (0.006) (0.149)
Complexity C Variance -34.101*** 3.834** 0.288* 0.476
(5.930) (1.423) (0.115) (0.687)
Constant 4.012*** 1.396*** 0.045*** 0.150*
(0.258) (0.200) (0.005) (0.067)
Obervations 11 11 11 11
Semirobust Standard errors in parentheses; *** p<0.001, ** p<0.01, * p<0.05 * + p<0.1
Table 1: ARMA predicting average similarity among visual arts in DeviantArt and Bēhance
Figure 3: Visualization of spatial relationships between the C-H space and the observed image diversity in DeviantArt and Bēhance data as measured by the image similarity. (a, b, c) Degrees of sample density, IE similarity (sample mean), and SIFT similarity (subsample mean) of the DeviantArt images per bin in the C-H region. (d, e, f) Degrees of sample density, IE similarity (sample mean), and SIFT similarity (subsample mean) of the Bēhance images per bin in the C-H region. Each C-H bin (C $\approx$ 0.01, H $\approx$ 0.02) per platform occupies more than 50 images.
The ARMA model is used separately for subsampled images from DeviantArt and Bēhance datatset. The results (Table 1) indicate that C and H values of a particular year are associated with an increase in diversity (over the previous year); the C and H values can affect the increase in diversity, despite the decrease in average absolute diversity over time. Here, we note that as the analysis predicts a movement of the average value of diversity for a given year with its average value of C and H, the area with large C and H values in the entire C-H space does not represent the area of high diversity. Meanwhile, the variance of H is added as an independent variable and controlled on the one hand, confirming that H variance also significantly has a positive correlation with diversity in all cases.
Higher H is associated with lesser degrees of IE and SIFT similarity among the artworks over time, as revealed by both platforms of our research. Similarly, higher levels of C are significantly associated with lesser degrees of IE and SIFT similarity. The H of artworks in both platforms are skewed to the left (DeviantArt: skewness = -1.95; Bēhance: skewness = -1.04), whereas the C is relatively symmetrical (DeviantArt: skewness = 0.2; Bēhance: skewness = 0.02).
A distribution of given years’ images in the C-H space is likely to include a wider variety of styles along with an increase in image diversity within the given year group upon the following conditions. First, when an average movement (direction of increases in C and H) occurs toward the upper right C-H region, which is the optimal direction of the upper left (similar, unexplored, sparse) and lower right (diverse, explored, dense) C-H region. Additionally, when the distribution (large H variance) is made on the entropy H axis.
We now partially examine the robustness of the ARMA results through patterns of the local stylistic diversity visualized over binned areas of a partitioned C-H region (C: 0 - 0.31, H: 0.5 – 1; see Fig. 3). Overall, the sample count plots (Fig. 3a, d) show that the most densely populated area is to the lower right C-H region with given embedding parameters, dx = dy = 2. This confirms that there is likely to be a strong correlation between image diversity and the highly concentrated C-H region, where previous art-historical styles with great image diversity are positioned. Hence, the more artworks where entropy H is greater, the greater image diversity there is likely to be.
In terms of observing image diversity manifested in the C-H region, we first measure the mean value of pairwise cosine similarity of IE of all the samples from each accountable C-H bin. Subsequently, we measure the mean value of pairwise Jaccard similarity of SIFT features of randomly subsampled artworks from each accountable C-H bin. The results altogether reveal that the highly dense C-H area tends to reflect less image similarity. The IE similarity (Fig. 3b, e) and SIFT similarity (Fig. 3c, f) plots demonstrate that at a given level of entropy H, higher complexity C is likely to be associated with greater intragroup image diversity (also see Fig. SI 4 for the supporting results of the WikiArt images previously mapped for Fig. 1a).
As observed in Fig. 1a, our results uncover the C-H movements of quasi-canonical visual artworks in contemporary online platforms, which ultimately trended to the upper-left C-H region. The C-H movements begin and propel away from the highly dense and stylistically diverse (low C and high H), and move towards sparser and stylistically homogenic areas (high C and low H), forming a process of a particular stylization. Such a stylization (towards higher C and lower H) is consistent with the broader movements in Western art history [13], which ranged from classical to modern times. However, the region with high C and low H has the lowest diversity (i.e., high image similarity) of image representations in artworks. Therefore, even though propelling towards the area is a new artistic endeavor in terms of C and H, there are few artworks that could be expressed with the given C-H values.
Based on our analyses, we suggest conditions for the groups of visual artworks to have high intragroup diversity, and shed light on an evolutionary process in which the intragroup diversity gives rise to styles. It is confirmed that when the average C-H movement of the groups are balanced towards higher C and higher H rather than the extremes in the C-H space, the stylistic diversity of the groups show a significant tendency to increase more compared to the previous period. Our findings also suggest that a visual artistic stylization process occurs as a result of the cumulative diversity of individualized artworks contributing as original mutations over time, to the transformative process of the prior narratives of styles, and emergence of a new style.
3 Discussion
Beyond computationally observing macroscopic evolution of visual art, opportunities for analytically mapping its latent dynamics in a creativity framework remain. Notably, Sigaki et al. [13] has performed a seminal work that demonstrates the feasibility of using the C-H plane to intuitively map the diachronic evolution of visual art styles. The study has recognized “a natural law in the same way as physical growth (Wölfflin)” when affiliating historical alterations in style, particularly from linear to painterly. This remark is similar to David Bohm’s belief about artistic creativity being related to “a harmony parallel to that of nature (Cézanne) [31].”
Apart from the findings of Sigaki et al.’s study providing useful insights into the evolution of visual art styles and its potential relationship to natural laws and creativity, we take an alternative strategy to explore the evolutionary dynamics of creative visual arts in the C-H plane. We look at the evolutionary process underpinning a stylization through the lens of selection, in which a variety of competing agents interact. This state arises by the emergence of novel and different mutations during the reproduction process [32]. A stylization of visual arts entails development of popularization of diverse artistic forms influenced by novel stylistic canons. Therefore, considering diversity as a prerequisite for the emergence of creativity and subsequent stylization, we investigate the C-H plane to empirically identify the optimal C-H condition, by which stylistic diversity is most pronounced while visual artworks of a specific timeframe and groups continue migrating towards homogenic styles.
We find that as the average entropy H rises, so does the intragroup diversity. Our analysis indicate that a novel style emerges in C-H regions where random and diverse agents coexist with strong individual innovativeness (Fig. 1). This finding echoes Eric Hobsbawm, who observed that during the period of highly unpredictable social and technological changes (i.e., the Avant Garde prior to 1914), modern art (low C and high H; Fig. 1a) was “not to claim that it displaced the classic and the fashionable, but that it supplemented both [3].”
The ARMA analysis (Table 1) supports our viewpoint by demonstrating a substantial increase in the respective group’s intragroup diversity when the group’s C-H movement is directed towards a balance (synchronous increases in C and H). Moreover, the results of this study empirically confirm that the intragroup image diversity varies depending on the locations in the C-H plane, further indicating the heterogeneity of image representations and local stylistic diversity of the C-H space. Therefore, from a collective rather than an individualist perspective, we assume that the optimal C-H condition for the greatest intragroup diversity and emergence of a certain style will be revealed when a group sets a balance between attempts to escape the conventionally dense C-H area (moving toward high C and low H areas) and attempts to remain in the C-H area where diversity can be expressed (low C and high H).
Possible limitations of our study include the possibility of selection bias in our data and methodologies for measuring image similarity. Although our dataset contains works by artists across the world, it is dominated by Western artists and is primarily composed of two-dimensional visual artworks. We are also aware that the curated user-generated visual artworks on DeviantArt and Bēhance are likely to represent merely a small portion of a colossal evolution occurring in contemporary visual art disciplines. Regarding our image similarity measures, despite the fact that the ResNet-18 architecture and SIFT capture multi-level image features, there are numerous other possible combinations of methods that capture various spatial image characteristics. Comparing diversity and creativity across different spaces of representation could be an intriguing future research topic.
Future research could investigate the properties of images generated by artificial intelligence (AI) models. In recent years, state-of-the-art AI models (e.g., DALL-E2 [33], stable diffusion models [34], etc.) have made enormous strides in text-to-image art generation. With growing interest in AI-generated images, the relationship between AI and human creativity has come to the forefront of research. A recent study of images generated by a stable diffusion model revealed that AI-generated images tend to have a narrower distribution of entropy and complexity when compared to human drawings; this may lead to less diversity and creativity in their visual artistic expression [35]. In addition, another study addressed the difficulties of evaluating interactions between human and computational creativity in online creative ecosystems [36]. Even though the development of AI art is rapidly advancing and attracting a great deal of attention, we cannot overlook the need to investigate and comprehend its characteristics in a more systematic manner. One could pose queries about the potentials of AI art, with a desire to investigate its representations and inspirations. As our methods are readily applicable to any corpus of two-dimensional digital images and their resulting knowledge maps, we believe the framework of this study can facilitate a better understanding of diversity and stylization in the emerging field of AI art.
4 Data (materials)
Quantitatively analyzing raw cultural data of user-generated content to capture their similarities on many possible dimensions can yield insights into the diversity of data’s visual organization in multidimensional spaces [37]. On this account, we assess popular online platforms as sources of artistic creativity and diversity - DeviantArt and Bēhance, which have become increasingly popular among artists who exhibit and share their creative works online. Accessibility, availability, and direct observation of a large corpus of user-generated visual art images are notable benefits of these platforms.
We use the 149,780 images with the C-H values (Table SI 1) during the IE obtaining process. However, the computing process of extraction and pairwise comparison of SIFT features between images consumes considerably more time compared to that of the image embeddings through convolutional neural networks (CNN). Therefore, we take two separate subsampling procedures for EDA on spatiotemporal relationships between local information of the C-H space and image diversity. 1) We randomly subsample images for both IE and SIFT similarity measures for the ARMA model (Table SI 3). 2) Considering each C-H bin sample as a population, we additionally use sample size formula (confidence level: 95%, margin of error: 5%) to determine and randomly subsample the minimum number of necessary samples to meet the desired statistical constraints, and to draw proper inferences from respective C-H bins in Fig. 3.
5 Methods
5.1 Complexity-entropy (C-H) measures of visual artwork images
Sigaki et al. [13] observed that the change in complexity (C) and entropy (H) of paintings over time could reflect the evolution of art historical styles. Here, we explain how a two-dimensional image is represented in the C-H space. The normalized permutation entropy H and statistical complexity C of an image are calculated based on its matrix representation. First, we investigate submatrices with ordinal patterns for each image with dx by dy embedding dimensions. The normalized Shannon entropy H(P) is then computed with the estimated probability distribution P, which is the probability of the order of neighboring pixel values in an image based on all ordinal patterns within each submatrix for a given embedding parameter (in our case, dx = dy = 2).
$H(P)=\frac{1}{\ln(n)}\sum_{i=0}^{n}p_{i}\ln\left(1/p_{i}\right)$ (1)
The value H signifies the degree of “disorder” in the configuration of the pixels of an image represented by its matrix representation. Moreover, the statistical complexity C(P) is calculated for investigating the degree of structural complexity present in each submatrix,
$C(P)=\frac{D(P,U)H(P)}{D^{*}}$ (2)
[…]
We note that in theory, any input image can be plotted as a point in the possible C-H domain bounded by a lower (minimum) and an upper (maximum) parabolic complexity curves, depending on the size of ordinal patterns (based on the embedded parameter (dx = dy = 2)) among the image pixels [38]. Specifically, the Jensen-Shannon divergence between a probability distribution P of local order patterns and the uniform distribution U forms a parabolic range as it is plotted against the given H value [13, 39, 40].
In practice, we calculated the C-H values of artworks using the Ordpy module [41], a simple and open-source Python module that implements permutation entropy and several principle methods for analyzing time series and two-dimensional data using complexity parameters [16, 42, 43].
5.2 Statistical regression of visual artwork image similarities
Among typical statistical models for time series analysis (e.g., the Bayesian network, the hidden Markov model, and etc.), an autoregressive moving-average (ARMA) model has been used to make inferences of relationships between variables by the form of regression. Parameters of the model are p (order of the autoregressive part specifying the number of lags used) and q (order of the moving average part representing the error of the model as a combination of previous error terms) [44].
We use the ARMA model to regress the variance of the similarities among the artworks on the degree of entropy H, complexity C, and their variances over the years of 2010 to 2020 (Table 1). The Modified Dickey-Fuller unit-root test is used to determine the number of lags required to achieve a stationary series [45], resulting in 1 lag for SIFT models, and 3 lags for IE models. Additionally, the mean and variance of the entropy H and complexity C are calculated using all the images from respective years.
5.3 Multi-level image features for visual artwork similarity measures
The Euclidean distance in the C-H space (Fig. 2a) is a straightforward metric for measuring the stylistic dissimilarity between images. In addition, we further adopt and utilize both global and local descriptors to extract multi-level image features and measure their similarity: a ResNet architecture and the SIFT algorithm. By utilizing the multi-level image features that aggregate both low- and high-level features, we disclose stylistic characteristics of image representations in the C-H plane.
Recent analyses of visual art styles use style vectors based on deep CNN to extract stylistic - i.e., high-level features and information from paintings [46]. On the other hand, the SIFT feature descriptor as a well-established CV technique has been demonstrated that the descriptor is effective for a variety of objectives related to image matching and object recognition. CNN filters by themselves perform similarly to SIFT in detecting low-level and straightforward - i.e., hand-crafted local invariant features, while outperforming SIFT in detecting high-level and complex features [47]. The combined usage of CNN and SIFT features has proven to result in discriminative multi-level image features using the best of both [48].
We initially build on ResNet-18 architecture pretrained with the ImageNet database to construct multi-level image features in our image set. ResNet pretrained with ImageNet has also been shown to provide sufficient feature representations for paintings [19], making it appropriate for extracting latent visual features - i.e. image embeddings (IE) - from heterogeneous visual artworks.
Inspired by Liu et al.’s approach of extracting multi-level features of artworks to represent art style of images [21], we obtain the embeddings of our images in the following manner (see Fig. 2b for the detailed technical pipeline). 1) As combining both low- and high-level features are effective for encoding art styles, the max (front) and average (back) pooling layers of ResNet-18 are chosen as the convolutional layers to extract hidden feature maps from each input image (224 × 224 × 3). 2) We then reduce dimensionality of the hidden low-level feature map (56 × 56 × 64) extracted from the maximum pooling layer to a one-dimensional feature map (56 × 56 × 1). 3) We also extract a hidden high-level feature map (1 × 512) from the average pooling layer just prior to the fully connected layer, where no feature map exists. 4) We then run a Principal Component Analysis (PCA) on both low-level and high-level feature maps to derive 100-dimensional embedding from each, thereby creating a 200-dimensional embedding vector of each user-generated artwork. 5) Lastly, we measure pairwise cosine similarities between embeddings of the contemporary user-generated visual artworks.
Following the neural network-based image similarity measurement, we utilize the SIFT feature matching technique (see Fig. 2c for the detailed technical pipeline). Specifically, SIFT uses the local histogram of oriented gradients to match the local features of an image with those of other images [49]. The SIFT features are extracted from keypoints detected between two distinct images while preserving their original data dimensions; gradient orientation histograms are extracted from quadrants of points of interest in an image. They are then merged into a normalized histogram (a SIFT descriptor) that is invariant to image location, scale, and rotation [20]. Each SIFT descriptor in an image is compared to its counterparts in other images.
We implement the standard SIFT feature indexing and matching technique for obtaining the desired degree of similarity between images based on its highest matching accuracy compared to the performance of other succeeding feature descriptors (e.g., PCA-SIFT, SURF, BRIEF, ORB, etc.) [50]. As the SIFT feature extractor, we use the xfeatures2d module from the cv2 interface in OpenCV versions for Python [51] to calculate the SIFT matching degree between images. 1) We first detect the local extrema in a single image as a potential keypoint detected in each pixel compared with its neighboring pixels. Each accurate keypoint selected from an image then generates an orientation histogram of 8 bins for the keypoints’ neighboring 16 × 16 blocks, which are then subdivided into 16 sub-blocks of size 4 × 4. The respective keypoints are subsequently allocated 128 (= 8 × 16) bin values as a vector [52]. 2) We then calculate the number of acceptable matches between keypoints from a pair of input images by the Fast Library for Approximate Nearest Neighbors (FLANN) based on a k-nearest neighbors (KNN) matcher [53]. 3) Using the number of keypoints and their good matches from each pair of images, we calculate the intersection (n of matching keypoints) divided by the union (n of total keypoints from a pair of images) of the images’ keypoints.
Data availability
The data underlying the analyses and findings of this study will be available from S.K. (ryankim990@gmail.com) on a reasonable request. Researchers interested in accessing data can contact the corresponding author (wnjlee@kaist.ac.kr). Correspondence and requests for data should be addressed to S.K. and W.L.
Competing interests
The authors declare no competing interests.